Darren is four times old as Tabitha. In twenty years, Darren will be twice as old as Tabitha. How old are they now?
Found 2 solutions by addingup, ankor@dixie-net.com:
Answer by addingup(3677)

D = 4T
and in twenty years:
D+20 = 2(T+20) substitute for D:
4T+20 = 2(T+20)
4T+20 = 2T+40
4T = 2T+20
2T = 20
T = 10
Tabitha is 10 and Darren is 40

Answer by ankor@dixie-net.com(22740)

Wrote an equation for each statement
:
Darren is four times old as Tabitha.
d = 4t
In twenty years, Darren will be twice as old as Tabitha.
d + 20 = 2(t+20)
d + 20 = 2t + 40
d = 2t + 40 - 20
d =2t + 20
replace d with 4t from the 1st statement
4t =2t + 20
4t - 2t = 20
2t = 20
t = 10 yrs is Tabitha's age now
then
4(10) = 40 yrs is Darren's age
:
:
:
Check solution in the statement
"In twenty years, Darren will be twice as old as Tabitha. "
40 + 20 = 2(10 + 20)